Grade 12Electrostatics

64 identical spheres each of charge q and capacitance c combine to form a large sphere. What is the capacitance of the larger sphere?

Archisman Bhunia
9 Years agoGrade 12

1 Answer

sharad
ApprovedApproved Tutor Answer9 Years ago
Capacitanve of solid sphere=4(pie)è×rCapacitance when combine =4(pie)èxr(new)As volume is constant therefore64×4/3(pie)r^3=4/3 (pie)R^3 R=4rTherefore capactinace of combine =4(the capacitance of one sphere)